A solid cylinder of cortical bone has a length of 500mm, diameter of 2cm and a Young’s Modulus of 17.4GPa. Determine the spring constant ‘k’
Answer #1

as we know

F = k dL

k = F /dL

also

dL = F L / A (Y)

A Y / L = F/dL

pi r^2 * Y / L = k

k = 3.14* 0.01^2* 17.4* 10^9 / 0.5

k = 1.093* 10^7 N/m

=======

comment before rate in case any doubt, will reply for sure.. goodluck